OpenStudy (anonymous):

come on this the 3rd time asking this is my last question please help please help If customer A has an appointment 2 2/3 hours before customer B who has an appointment 5/6 of an hour before customer C. How much later is customer C‘s appointment than customer A’s?

OpenStudy (anonymous):

3 1/2 hours.

OpenStudy (radar):

This is an addition problem 2 2/3 + 5/6 =16/6 + 5/6=21/6=3 1/2 hrs